OpenStudy (anonymous):

what is the slope between the two points (5, 1) and (4, 6)? My answer is 1/5 Am I right? :)

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

and no you are wrong

OpenStudy (johnweldon1993):

Use the equation @completeidiot just put above lol \[m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] You have the points (5, 1) and (4, 6)? (x1,y1) (x2,y2) so plug those in and what do you get?

OpenStudy (anonymous):

the answer would be 5/-1

OpenStudy (anonymous):

now simplify it

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

you dont have to simplify it but it would be 5/-1 or -5

OpenStudy (johnweldon1993):

That would be correct @hunterhoff

OpenStudy (anonymous):

depending on the teacher you have

OpenStudy (anonymous):

Thank you guys! :D

OpenStudy (anonymous):

your welcome